OpenStudy (anonymous):

The 12th term of an arithmetic sequence is 36, and the common difference is 3. What is the first term of the sequence? A. 1 B. 2 C. 3 D. 4

OpenStudy (anonymous):

nth term is given as\[a_{n} = a + (n-1)d\] where a = first term , n = no. of terms, d = common difference

OpenStudy (cwrw238):

an arithmetic sequence is a, a+d, a+2d, a + 3d +... where a = first term and d = common difference

OpenStudy (cwrw238):

so the 12th term = a + 11d and this = 36 giving a + 11*3 = 36 find a

OpenStudy (anonymous):

ok I got it thanks CWRW & your pic <3 lol